OpenStudy (anonymous):

Volume of the region bounded by y=e^(x-4), y=0, x=4, x=5. Rotated around x-axis.

OpenStudy (anonymous):

I like these problems, but I like it even more when @ganeshie8 explains it :)

OpenStudy (neer2890):

Volume of revolution about x-axis of the area bounded by curve y=f(x) , the x-axis and the ordinates x=a,x=b is\[\int\limits_{a}^{b} \pi y ^{2}dx\]

OpenStudy (neer2890):

so here y=e^(x-4),a=4,b=5 so \[volume=\int\limits_{4}^{5}\pi [e ^{(x-4)}]^{2} dx\]

OpenStudy (neer2890):

now you can easily find the volume by integrating the above equation.

OpenStudy (anonymous):

Thanks! :D All the bounds and limits threw me off =/

OpenStudy (neer2890):

you're welcome...:)